# The dawn is quiet II: Gaia XP constraints on the Milky Way’s proto-Galaxy from very metal-poor MDF tails

### v1

@@ 2 Gaia XP metallicity catalogs
3 Galactic chemical evolution model
We use the same 735 GCE models that successfully replicated the moderate rise in [Mg/Fe] as seen in the H3 survey from our Paper I. Our model framework is designed to investigate the early $\alpha$/Fe] behavior in the proto-Milky Way, simulating the first two Gyr of its evolution within a bin with 1 kpc radius. The model divides this period into two one-Gyr phases, with time steps of 30 Myr, reflecting the longest core-collapse supernovae (CCSNe) progenitor lifespans. Stellar populations, ranging from 0.5 to 40 M and following an initial mass function (IMF) of $\xi(m)\propto m^{-2.3}$, are tracked alongside cold and warm interstellar medium (ISM) components. Nucleosynthetic yields from 2013ARA&A..51..457N are interpolated across metallicities to determine the chemical enrichment from stars. Stars between 0.5 and 0.9 M are formed and store cold ISM, but they don’t contribute to nucleosynthesis due to the lack of nucleosynthesis yields.
[…]
4 Results
@@
Figure 2: Stellar distribution in Galactocentric rotational velocity versus metallicity for various Gaia XP metallicity catalogs. The red dashed line in every panel indicates the median velocity across the metallicity range. Most metal-poor stars have low rotational velocities as expected for the proto-Galaxy, but false metal-poor stars that likely belong to the disk become common as [Fe/H] exceeds -2.0. Similar contamination also appears in 2024MNRAS.531.2126F and 2025ApJS..279....7Y as [Fe/H] drops below -2.6.
Figure 4: Gaia CMD colored by median Galactocentric rotational velocity ($v_{\phi}$) in various Gaia XP metallicity catalogs. The magnitudes are observed absolute magnitudes as in b) in Fig. 1. The slow-rotating stars highlight a PARSEC 1.2S isochrone for a stellar population 13 Gyr old and with [M/H] = -$2.$ Most stars with radial velocities in the various catalogs also tend to reside along such an isochrone on the CMD as seen in Figure 1. This suggests that at least most stars with low $v_{\phi}$ in Gaia RVS are genuinely old metal-poor stars.
4.1 Zero Rotation Expected for the Proto-Galaxy
Figure 1 shows the stellar density distribution in the Gaia color-magnitude diagram (CMD) for various Gaia XP spectroscopic metallicity catalogs. The top two rows (panels (a)) use apparent magnitudes $G$ and BP-RP colors, while the bottom two rows (panels (b)) use observed absolute magnitudes $G_{\mathrm{obs}}$ derived from parallax without extinction or reddening corrections.
@@ 4.1 Zero Rotation Expected for the Proto-Galaxy
Panels a) provide a visual representation of the quality cuts per the authors’ recommendations listed in the last column of Table 1, while panels (b) reduce distance-driven smearing and makes the CMDs more comparable across catalogs, at the cost of relying on parallax quality. We do not perform dereddening because the working samples have low extinction after the quality cuts.
@@
Stars with [Fe/H]<1.5<-1.5 are expected to have near-zero mean rotation with few objects reaching disk-like azimuthal velocities $v_{\phi}\gtrsim 100~\mathrm{km\,s^{-1}}$ (2000AJ....119.2843C; 2007Natur.450.1020C; 2011MNRAS.411.1480D; 2018ApJ...856L..26M; 2024A&A...691L...1Y). To test this expectation and diagnose catalog-dependent artifacts, we derived Galactocentric coordinates ($R$, $\phi$, $z$) and velocities ($v_{r}$, $v_{\phi}$, $v_{z}$) from Gaia astrometry using astropy (2022ApJ...935..167A). Figure 2 presents the two-dimensional stellar density distribution of $v_{\phi}$ versus [Fe/H] for $3\leq[\mathrm{Fe/H}]<-1.5$ in the metallicity catalogs. The running median $v_{\phi}$ is marked in each panel with a red dashed line. The catalog by 2024arXiv240900197K applied a cut of $v_{\phi}<160$ $\mathrm{km\,s^{-1}}$.
@@
In several catalogs, we observe a concentration of stars with $v_{\phi}\approx 220$ $\mathrm{km\,s^{-1}}$ as [Fe/H] approaches our boundary (typically $1.8\lesssim[\mathrm{Fe/H}]<-1.5$), most prominently in 2023MNRAS.524.1855Z; 2024MNRAS.531.2126F; 2024A&A...692A.115M; 2025ApJS..279....7Y. At the very metal-poor end with [Fe/H] <2.5<-2.5, the median $v_{\phi}$ differs across catalogs: in 2024MNRAS.531.2126F; 2025A&A...695A..75Y; 2025ApJS..279....7Y it exceeds $50~\mathrm{km\,s^{-1}}$, while in the remaining catalogs it stays closer to 0. The high-$v_{\phi}$ features near the metallicity boundaries can arise from: (1) misclassification of metal-rich disk stars as metal-poor; (2) genuine metal-poor stars on fast, disk-like orbits associated with the metal-weak thick disk or ancient, rotation-dominated components (2021MNRAS.508.1509F; 2024A&A...683A.136B; 2024A&A...691L...1Y; 2025NatAs...9..101X).
@@
A physically motivated channel for genuinely fast, metal-poor stars could be an early, gas-rich merger that seeded the proto-disk’s angular-momentum axis, producing a prograde, low-$\mathrm{[Fe/H]}$ tail (2021MNRAS.505..921S). An alternative is that later mergers, such as the GSE, could have torqued pre-existing stellar populations, imparting net rotation (2024MNRAS.527.6926M). Whether the observed prograde motions are primordial or merger-induced, their kinematic coherence suggests a link to the early building blocks of the proto-Galaxy (2019MNRAS.484.2166S; 2021ApJ...908..208C). However, as shown in Fig. 2, the bulk of stars with $\mathrm{[Fe/H]}<-2.0$ exhibit low $v_{\phi}$, consistent with a dispersion-dominated proto-Galaxy. Because only $\sim 5$–10% of stars in each catalog have Gaia RVS radial velocities, kinematic cuts alone cannot be applied to the entire catalogs. An alternative approach is to apply isochrone-based selections, which can be done with the available photometry and helps mitigate contamination. We restrict our MDF analysis to $\mathrm{[Fe/H]}<-2.0$ to minimize the inclusion of disk stars. Since most stars already follow the observed isochrone for an old, metal-poor stellar population and most RVS stars have low rotation, we leave a detailed kinematic decomposition for future work.
@@
Figure 3 shows the two-dimensional number density in Galactocentric radius $R$ and height $|\mathrm{z}|$ for the XP catalogs we used within our metallicity range. Unlike many previous studies that targeted high-$|\mathrm{z}|$ regions, our working samples are overwhelmingly dominated by stars in the disk plane: most stars lie at $|\mathrm{z}|<2$ kpc because the underlying Gaia XP selections are magnitude-limited and largely all-sky. The vetted RGB catalog by 2023ApJS..267....8A and the catalogs by 2025MNRAS.536.2507K and 2025ApJS..279....7Y target giants, which results in more even $R$ coverage but naturally misses the dwarfs at low-$|\mathrm{z}|$. The pipeline by 2023ApJS..267....8A require WISE $W1/W2$ and even use $M_{W1}$ and GWG-W cuts to purify giants. CatWISE was adopted precisely to improve completeness over AllWISE in crowded fields. Consequently, some residual suppression in the dustiest mid-plane lines of sight is expected.
@@
Figure 4 maps median $v_{\phi}$ across the CMD (using $G_{\mathrm{obs}}$), localizing, via the colour scale spanning from 100-100 to $300~\mathrm{km\ s^{-1}}$, where prograde stars concentrate in evolutionary phase. Stars with median $|v_{\phi}|<50~\mathrm{km\ s^{-1}}$ trace a PARSEC 1.2S isochrone for an old, metal-poor population ($\mathrm{[M/H]}=-2.2$, $\mathrm{Age}=13\,\mathrm{Gyr}$), as expected, while stars that deviate from the isochrone—typically toward the blue turnoff and red subgiant/dwarf edge—show elevated $v_{\phi}$ over $200~\mathrm{km\ s^{-1}}$. These concentrations with high $v_{\phi}$ on the CMD coincide with regions where training set coverage is sparse because more than half the catalogs are trained on or anchored to APOGEE. For example, 2024MNRAS.531.2126F trained their NNs on APOGEE data and 2025A&A...695A..75Y used APOGEE to train their NNs for flux correction. Even broader “transfer” models like SHBoost by 2024A&A...691A..98K still draw much of their label scale from spectroscopic surveys that under-sample parts of the CMD and thus inherit the coverage gaps. Therefore, the high-$v_{\phi}$ patches are most naturally interpreted as thin-disk contaminants whose metallicities are severely underestimated by the pipelines.
@@
A similar trend is seen in independent photometric-metallicity work by 2024A&A...692A.115M: CaHK-syn metallicity estimates degrade toward hotter stars and lose sensitivity for dwarfs. Substituting their default (giant-calibrated) metallicities with the companion dwarf-model values did not alter the predominance of high-$v_{\phi}$ dwarfs. The difference in [Fe/H] between the two models is generally 0.2–0.4 dex, which is small compared to our metallicity range. Catalogs that confine their labels to well-calibrated regions of the CMD generally avoid spurious high-$v_{\phi}$ patches. For example, 2023ApJS..267....8A released a vetted RGB-only catalog. 2025MNRAS.537.1984A also selected their objects partially from this RGB sample. More broadly, a training-set-dependent CMD mask-such as the flag provided by 2025ApJ...980...90H-offers a practical way to refine samples and minimize contamination.
@@
Figure 5: Logarithmic metal-poor MDFs from various Gaia XP catalogs after applying the quality cuts listed in Table 1. Dashed lines show exponential density profiles with reference slopes $k=0.6$, 1.0, and 1.4 for comparison. Despite differences among the catalogs, most exhibit a near-exponential profile with $k\approx 1.0$. Catalogs from 2024A&A...691A..98K and 2025A&A...695A..75Y, trained solely on APOGEE, under-represent VMP stars.
@@
Figure 6: Logarithmic metal-poor MDFs from Gaia XP metallicity catalogs as a function of Galactocentric radius ($R$) and observed magnitude ($G_{\mathrm{obs}}$). Panels in a) show MDFs in bins of $R$, incremented by 2 kpc from 0 to 12 kpc. Panels in b) show MDFs in bins of $G_{\mathrm{obs}}$. In all panels, dashed lines indicate exponential density profiles with three reference slopes ($k=0.6,1.0,1.4$) for comparison as in Fig. 5. The MDF slopes are consistent with selection by $R$ or $G_{\mathrm{obs}}$, except when there are too few stars to form MDFs.

### v2

@@ II.2 Training Methods for Gaia Catalogs
Because the Gaia XP metallicity catalogs do not share a common selection function, we use the published metallicities primarily for differential analyses. We do not attempt to construct a volume-complete sample or combine all catalogs into a single homogenized dataset, but instead treat each catalog within its published quality-controlled regime. Several potential biases are relevant. First, metal-poor giants are intrinsically brighter than dwarfs, so magnitude-limited samples can over-represent giants and probe different Galactic volumes and component mixtures depending on the CMD coverage of each catalog. Second, because the Gaia XP samples span both disk and halo sightlines, contamination by metal-rich disk stars with underestimated metallicities may affect the metal-richer boundary of the inferred metal-poor tail. Since these effects depend on the training set, inference domain, and recommended quality cuts of each public release, they must be evaluated empirically when interpreting the MDFs (see Section 4.2).
@@
III Galactic chemical evolution model
We use the same 735 GCE models that successfully replicated the moderate rise in [Mg/Fe] as seen in the H3 survey (Conroy et al., 2019) from our Paper I. Our model framework is designed to investigate the early $\alpha$/Fe] behavior in the proto-Milky Way, simulating the first two Gyr of its evolution within a bin with 1 kpc radius. The model divides this period into two one-Gyr phases, with time steps of 30 Myr, reflecting the longest core-collapse supernovae (CCSNe) progenitor lifespans. Stellar populations, ranging from 0.5 to 40 M and following an initial mass function (IMF) of $\xi(m)\propto m^{-2.3}$, are tracked alongside cold and warm interstellar medium (ISM) components. Nucleosynthetic yields from Nomoto et al. (2013) are interpolated across metallicities to determine the chemical enrichment from stars. Stars between 0.5 and 0.9 M are formed and store cold ISM, but they don’t contribute to nucleosynthesis due to the lack of nucleosynthesis yields.
The model incorporates key astrophysical processes. Star formation is governed by the Kennicutt-Schmidt law, converting cold ISM into new stellar populations. Stellar evolution dictates the return of enriched gas: CCSNe from stars more massive than 9 M explode within one time step (30 Myr), while intermediate-mass stars (3–8 M) evolve into white dwarfs (WDs) that can later produce Type Ia supernovae (SNe Ia). The SNe Ia rate depends on the WD mass reservoir, a constant fraction ($f_{\text{SNIa}}$), and an exponential delay timescale. Enriched material ejected from stars is distributed primarily to the warm ISM (79%) and a smaller fraction to the cold ISM (1%), with rest lost to the circum-galactic medium (CGM). The warm ISM cools to replenish the cold ISM, while feedback from star formation and supernovae heats a portion of the cold ISM into the warm ISM, regulated by mass-loading factors, $\eta_{\mathrm{SF}}=2.0$ and $\eta_{\mathrm{SN}}=5.0$. The outflow mass-loading factors are fixed and directly proportional to the mass involved in star formation and stellar evolution activities. The model also allows for pristine gas inflow into the warm ISM.
To understand the conditions leading to an early $\alpha$/Fe] rise, we explore a range of five fundamental free parameters: the initial cold ISM mass ($m_{0,\text{cold}}$), the SNe Ia fraction ($f_{\text{SNIa}}$), the warm ISM cooling timescale ($t_{\text{cool}}$), the star formation efficiency ($\epsilon_{\text{SF}}$), and the gas inflow rate ($\dot{m}_{\text{inflow}}$). These parameters are varied across astrophysically motivated ranges in numerous model runs. Notably, to achieve an $\alpha$/Fe] rise after an initial decline (due to SNe Ia), the star formation efficiency and inflow rate, along with inflow gas composition, are allowed to change in the second Gyr of the simulation, facilitating increased $\alpha$-element production through enhanced CCSNe activity or accretion of $\alpha$-enriched gas.
IV Results
@@
IV.1 Zero Rotation Expected for the Proto-Galaxy
Figure 1 shows the stellar density distribution in the Gaia color-magnitude diagram (CMD) for various Gaia XP spectroscopic metallicity catalogs. Panels (a) use apparent magnitudes $G$ and BP-RP colors, while the panels (b) use observed absolute magnitudes $G_{\mathrm{obs}}$ derived from parallax without extinction or reddening corrections.
@@ IV.1 Zero Rotation Expected for the Proto-Galaxy
Panels a) provide a visual representation of the quality cuts per the authors’ recommendations listed in the last column of Table 1, while panels (b) reduce distance-driven smearing and makes the CMDs more comparable across catalogs, at the cost of relying on parallax quality. We do not apply an additional dereddening correction because the retained samples have already been filtered into relatively low- or moderate-extinction regimes by the adopted catalog-specific quality cuts (E(BV)<0.1E(B-V)<0.10.50.5), and several pipelines explicitly account for extinction or operate on corrected XP spectra (see Table 1).
@@
Figure 2: Stellar distribution in Galactocentric rotational velocity versus metallicity for various Gaia XP metallicity catalogs. The solid red curve shows the running median $v_{\phi}$ as a function of metallicity, computed in a sliding window of width 0.2 dex in $\mathrm{[Fe/H]}$. The dotted lines show the 16th and 84th percentile range of $v_{\phi}$ respectively within the same window. Most metal-poor stars have low rotational velocities as expected for the proto-Galaxy, but false metal-poor stars that likely belong to the disk become common as [Fe/H] exceeds -2.0. Similar contamination also appears in Fallows & Sanders (2024) and Yang et al. (2025) as [Fe/H] drops below -2.6.
Figure 4: Gaia CMD colored by median Galactocentric rotational velocity ($v_{\phi}$) in various Gaia XP metallicity catalogs. The magnitudes are observed absolute magnitudes as in b) in Fig. 1. The slow-rotating stars highlight a PARSEC 1.2S isochrone (Bressan et al., 2012) for a stellar population 13 Gyr old and with [M/H] = -$2.$ Most stars with radial velocities in the various catalogs also tend to reside along such an isochrone on the CMD as seen in Figure 1. This suggests that at least most stars with low $v_{\phi}$ in Gaia RVS are genuinely old metal-poor stars.
To assess possible contamination, we derived Galactocentric positions and velocities only for the subset of stars with published Gaia DR3 RVS radial velocities, which constitutes approximately 5–10% of each Gaia XP metallicity catalog after the adopted quality cuts. These kinematic data are used solely as an external diagnostic of the metal-poor population and were not used to define the MDF samples or to impose kinematic cuts on the full catalogs. Stars with [Fe/H]<1.5<-1.5 are expected to have near-zero mean rotation with few objects reaching disk-like azimuthal velocities $v_{\phi}\gtrsim 100~\mathrm{km\,s^{-1}}$ (Chiba & Beers, 2000; Carollo et al., 2007; Deason et al., 2011; Myeong et al., 2018; Yuan et al., 2024). To test this expectation and diagnose catalog-dependent artifacts, we transformed the Gaia astrometry together with the available Gaia DR3 RVS radial velocities for this subset into Galactocentric coordinates ($R$, $\phi$, $z$) and velocities ($v_{r}$, $v_{\phi}$, $v_{z}$) using astropy (Astropy Collaboration et al., 2022). Because Gaia RVS radial velocities are available for only a minority of XP sources, these kinematic distributions should be interpreted as a validation subset rather than as a complete description of the parent metallicity catalogs.
@@
Figure 2 presents the two-dimensional stellar density distribution of $v_{\phi}$ versus [Fe/H] for $3\leq[\mathrm{Fe/H}]<-1.5$ in the metallicity catalogs. The running median $v_{\phi}$ is marked in each panel with a red solid line and the 16th and 84th percentile with dotted lines. The catalog by Kane et al. (2024) applied a cut of $v_{\phi}<160$ $\mathrm{km\,s^{-1}}$. In several catalogs, we observe a concentration of stars with $v_{\phi}\approx 220$ $\mathrm{km\,s^{-1}}$ as [Fe/H] approaches our boundary (typically $1.8\lesssim[\mathrm{Fe/H}]<-1.5$), most prominently in Zhang et al. (2023); Fallows & Sanders (2024); Martin et al. (2024); Yang et al. (2025). At the very metal-poor end with [Fe/H] <2.5<-2.5, the median $v_{\phi}$ differs across catalogs: in Fallows & Sanders (2024); Ye et al. (2025); Yang et al. (2025) it exceeds $50~\mathrm{km\,s^{-1}}$, while in the remaining catalogs it stays closer to 0. The high-$v_{\phi}$ features near the metallicity boundaries can arise from: (1) misclassification of metal-rich disk stars as metal-poor; (2) genuine metal-poor stars on fast, disk-like orbits associated with the metal-weak thick disk or ancient, rotation-dominated components (Fernández-Alvar et al., 2021; Bellazzini et al., 2024; Yuan et al., 2024; Xiang et al., 2025).
@@
A physically motivated channel for genuinely fast, metal-poor stars could be an early, gas-rich merger that seeded the proto-disk’s angular-momentum axis, producing a prograde, low-$\mathrm{[Fe/H]}$ tail (Santistevan et al., 2021). An alternative is that later mergers, such as the GSE, could have torqued pre-existing stellar populations, imparting net rotation (McCluskey et al., 2024). Whether the observed prograde motions are primordial or merger-induced, their kinematic coherence suggests a link to the early building blocks of the proto-Galaxy (Sestito et al., 2019; Carter et al., 2021). However, as shown in Fig. 2, the bulk of stars with $\mathrm{[Fe/H]}<-2.0$ exhibit low $v_{\phi}$, consistent with a dispersion-dominated proto-Galaxy. Because only $\sim 5$–10% of stars in each catalog have Gaia RVS radial velocities, kinematic cuts alone cannot be applied to the entire catalogs. An alternative approach is to apply isochrone-based selections, which can be done with the available photometry and helps mitigate contamination. We restrict our MDF analysis to $\mathrm{[Fe/H]}<-2.0$ to minimize the inclusion of disk stars. Since most stars already follow the observed isochrone for an old, metal-poor stellar population and most RVS stars have low rotation, we leave a detailed kinematic decomposition for future work.
@@
Figure 3 shows the two-dimensional number density in Galactocentric radius $R$ and height $|\mathrm{z}|$ for the XP catalogs we used within our metallicity range. Unlike many previous studies that targeted high-$|\mathrm{z}|$ regions, our working samples are overwhelmingly dominated by stars in the disk plane: most stars lie at $|\mathrm{z}|<2$ kpc because the underlying Gaia XP selections are magnitude-limited and largely all-sky. The vetted RGB catalog by Andrae et al. (2023a) and the catalogs by Kane et al. (2025) and Yang et al. (2025) target giants, which results in more even $R$ coverage but naturally misses the dwarfs at low-$|\mathrm{z}|$. The pipeline by Andrae et al. (2023a) require WISE $W1/W2$ and even use $M_{W1}$ and GWG-W cuts to purify giants. CatWISE was adopted precisely to improve completeness over AllWISE [^2] https://irsa.ipac.caltech.edu/data/WISE/docs/release/AllWISE/ in crowded fields. Consequently, some residual suppression in the dustiest mid-plane lines of sight is expected.
@@
Figure 4 maps median $v_{\phi}$ across the CMD (using $G_{\mathrm{obs}}$), localizing, via the colour scale spanning from 100-100 to $300~\mathrm{km\ s^{-1}}$, where prograde stars concentrate in evolutionary phase. Stars with median $|v_{\phi}|<50~\mathrm{km\ s^{-1}}$ trace a PARSEC 1.2S isochrone for an old, metal-poor population ($\mathrm{[M/H]}=-2.2$, $\mathrm{Age}=13\,\mathrm{Gyr}$), as expected, while stars that deviate from the isochrone—typically toward the blue turnoff and red subgiant/dwarf edge—show elevated $v_{\phi}$ over $200~\mathrm{km\ s^{-1}}$. These concentrations with high $v_{\phi}$ on the CMD coincide with regions where training set coverage is sparse because more than half the catalogs are trained on or anchored to APOGEE. For example, Fallows & Sanders (2024) trained their NNs on APOGEE data and Ye et al. (2025) used APOGEE to train their NNs for flux correction. Even broader “transfer” models like SHBoost by Khalatyan et al. (2024) still draw much of their label scale from spectroscopic surveys that under-sample parts of the CMD and thus inherit the coverage gaps. Therefore, the high-$v_{\phi}$ patches are most naturally interpreted as thin-disk contaminants whose metallicities are severely underestimated by the pipelines.
@@
A similar trend is seen in independent photometric-metallicity work by Martin et al. (2024): CaHK-syn metallicity estimates degrade toward hotter stars and lose sensitivity for dwarfs. Substituting their default (giant-calibrated) metallicities with the companion dwarf-model values did not alter the predominance of high-$v_{\phi}$ dwarfs. The difference in [Fe/H] between the two models is generally 0.2–0.4 dex, which is small compared to our metallicity range. Catalogs that confine their labels to well-calibrated regions of the CMD generally avoid spurious high-$v_{\phi}$ patches. For example, Andrae et al. (2023a) released a vetted RGB-only catalog. Ardern-Arentsen et al. (2025) also selected their objects partially from this RGB sample. More broadly, a training-set-dependent CMD mask-such as the flag provided by Hattori (2025)-offers a practical way to refine samples and minimize contamination.
IV.2 Robust MDF Slope Across Catalogs and Selections
@@
Figure 5: Logarithmic metal-poor MDFs from various Gaia XP catalogs after applying the quality cuts listed in Table 1. Bin width is for this and the panels in the following two figures is 0.02 dex. Dashed lines show exponential density profiles with reference slopes $k=0.6$, 1.0, and 1.4 for comparison. Despite differences among the catalogs, most exhibit a near-exponential profile with $k\approx 1.0$. Catalogs from Khalatyan et al. (2024) and Ye et al. (2025), trained solely on APOGEE, under-represent VMP stars. Although this and the following figures could extend to $\mathrm{[Fe/H]}>-2.0$, all slope estimates are restricted to $3<\mathrm{[Fe/H]}<-2.0$.
@@
Figure 6: Logarithmic metal-poor MDFs from Gaia XP metallicity catalogs in bins of Galactocentric radius $R$. The panels show MDFs in 2 kpc intervals from 0 to $12$ kpc. In each panel, dashed lines indicate exponential density profiles with three reference slopes ($k=0.6,1.0,1.4$) for comparison, as in Fig. 5. The inferred MDF slopes remain broadly consistent across radial bins, except where the number of stars is too small to define the MDF robustly.
Figure 7: Logarithmic metal-poor MDFs from Gaia XP metallicity catalogs in bins of parallax-corrected magnitude, $G_{\mathrm{obs}}$. The panels show MDFs for subsamples divided by $G_{\mathrm{obs}}$ estimated from the observed $G$ magnitude and parallax, without correcting for extinction. In each panel, dashed lines indicate exponential density profiles with three reference slopes ($k=0.6,1.0,1.4$) for comparison, as in Fig. 5. The inferred MDF slopes remain broadly consistent across $G_{\mathrm{obs}}$ bins, except where the number of stars is too small to define the MDF robustly.
To measure the slope of the metal-poor tail, we fit each catalog over the fixed interval $3\leq\mathrm{[Fe/H]}\leq-2$. Rather than fitting a straight line to binned histograms, we use the individual stellar metallicities directly. This avoids sensitivity to bin size and reduces the impact of noisy low-count bins.
We model the MDF tail as an exponential function,
$p(x)\propto\exp(\beta x),\qquad x\equiv\mathrm{[Fe/H]},$ (2)
which is equivalent to a straight line in logarithmic space,
$\log_{10}p(x)=kx+c,\qquad k=\beta/\ln 10.$ (3)
We determine $k$ with a maximum-likelihood fit using only stars in the chosen metallicity range. To estimate the uncertainty, we use bootstrap resampling. For each catalog, we repeatedly resample the stars with replacement, refit the slope, and take the median and 16th–84th percentile range of the resulting slope distribution. We report these as the best-fit value and uncertainty for each catalog in Table 2.
Table 2: MDF tail slopes estimated over $3\leq\mathrm{[Fe/H]}\leq-2$ using an unbinned maximum-likelihood fit to a truncated exponential model. Reported values are the bootstrap median and 16th–84th percentile interval.
Color Catalog $k$
  Andrae et al. (2023) $1.93^{+0.01}_{-0.01}$
  Zhang et al. (2023) $2.09^{+0.00}_{-0.00}$
  Khalatyan et al. (2024) $3.44^{+0.03}_{-0.03}$
  Hattori (2024) $0.92^{+0.00}_{-0.00}$
  Fallows & Sanders (2024) $1.65^{+0.01}_{-0.01}$
  Martin et al. (2024) $0.79^{+0.00}_{-0.00}$
  Kane et al. (2024) $1.40^{+0.02}_{-0.02}$
  Ardern-Arentsen et al. (2025) $1.26^{+0.02}_{-0.02}$
  Ye et al. (2025) $2.07^{+0.01}_{-0.01}$
  Yang et al. (2025) $0.63^{+0.00}_{-0.00}$
Figure 5 shows the MDFs over $3<\mathrm{[Fe/H]}<-1.8$ on a logarithmic scale for the Gaia XP catalogs listed in Sec. II. Using log density highlights the near-exponential falloff of the metal-poor tail, as in Fig. 4 of Rix et al. (2022) for $2<\mathrm{[Fe/H]}<-1$. For visual guidance, dashed lines indicate three reference slopes ($k=0.6,1.0,1.4$) that bracket the range we observe across catalogs in this interval. We excluded Li et al. (2024), whose metallicities are truncated at $\mathrm{[Fe/H]}=-2.0$, producing an artificial pile-up at the boundary.
Because Khalatyan et al. (2024) and Ye et al. (2025) are anchored to APOGEE training labels, their MDFs decline rapidly below $\mathrm{[Fe/H]}\approx-2$, where APOGEE offers few measurements. In addition, Ye et al. (2025) disable their APOGEE-based NN flux correction for stars with $\mathrm{[Fe/H]}\leq-2.5$ and fit FERRE directly, which they release as a separate metal-poor subset. This treatment imprints a discontinuity near $\mathrm{[Fe/H]}=-2.5$, but the MDF tail slope remains consistent. Yang et al. (2025) train against metal-poor reference sets (PASTEL, SAGA) with a cost-sensitive objective to enhance recovery of VMP giants, similarly flattening their tail. Finally, Martin et al. (2024)-as part of the Pristine effort targeting $\mathrm{[Fe/H]}<-1$-deliver a comparatively larger fraction of metal-poor stars from synthetic CaHK, which also yields a flat density decline in their MDF.
The similarly flat MDF tail in the Hattori (2025) catalog is best explained by two design choices that boost recovery of VMP halo stars: (1) the training set was deliberately augmented with the VMP sample of Li et al. (2022), extending label coverage well below $\mathrm{[M/H]}\approx-2.5$ and increasing counts at the low-metallicity end; and (2) the recommended catalog is restricted to low-extinction sight lines, E(BV)<0.1E(B-V)<0.1, which preferentially probe high-latitude halo volume and improves completeness for distant halo giants. However, the same VMP set was also used for training by Zhang et al. (2023) and Ardern-Arentsen et al. (2025), yet both deliver noticeably steeper metal-poor tails. This suggests that simply augmenting training set with VMP stars does not necessarily determine the slope; sample selection and quality cuts play a more important role.
Andrae et al. (2023a) and Zhang et al. (2023) appear to under-represent the most metal-poor regime: their MDFs show the steepest slope for $\mathrm{[Fe/H]}<-2.0$ with $k>1.4$, then flatten toward $k\approx 1.0$ as metallicity increases. Andrae et al. (2023a) selected sources in their vetted RGB sample to avoid spurious metal-poor stars, which could preferentially remove some VMP stars. Zhang et al. (2023) did not explicitly include VMP stars in their training set and their basic reliability cuts tend to cull low-SNR sources, disproportionately trimming the VMP tail. Overall, the metal-poor tails have logarithmic slopes centered near $k\approx 1$, and are well described by a single exponential over our metallicity window. We observe no bumps or peaks within $3<\mathrm{[Fe/H]}<-1.8$ after applying recommended quality cuts from the authors. Minor departures from linearity in log-density appear only at the boundaries-flattening in VMP-boosted catalogs and steepening where training or completeness drops below $\mathrm{[Fe/H]}\approx-2.0$-and are consistent with known selection and pipeline effects.
Figure 6 shows MDFs over $3<\mathrm{[Fe/H]}<-1.8$ split by Galactocentric radius $R$, while Figure 7 shows the same MDFs split by parallax-corrected magnitude, with the same reference slopes from Fig. 5. Because the $y$-axis is logarithmic, the curves tend to meet near $\mathrm{[Fe/H]}\simeq-1.8$ at density of 1. A quick way to read an individual slope is to note the intercept at $\mathrm{[Fe/H]}=-3$: for example, an intercept of $10^{-1}$ implies $k\approx 1.0/1.5\simeq 0.67$, while $10^{-2}$ implies $k\simeq 1.33$. Most catalogs exhibit nearly identical slopes across the various $R$ and parallax-corrected magnitude bins. The consistency across radius indicates that the MDF slope is not strongly sensitive to distance from the Galactic center, while the consistency across parallax-corrected magnitude suggests that the result is robust across different stellar types and luminosity classes despite the differing CMD footprints of the catalogs.
In Fig. 7, several samples contain fewer VMP stars at fainter parallax-corrected magnitudes. This is expected because some catalogs are dominated by giants and therefore contain relatively few intrinsically faint stars, while others have smaller overall samples in those bins. The MDF by Zhang et al. (2023) loses VMP stars more rapidly toward the faint end, consistent with their basic reliability cut, which preferentially culls lower-S/N sources with larger uncertainties. By contrast, the catalogs from Hattori (2025), Yang et al. (2025), and Martin et al. (2024) show nearly constant slopes across the parallax-corrected magnitude range. The first catalog supplemented its training set with VMP stars and uses a CMD-based quality filter that retains MSTO and main-sequence stars, while the latter two benefit from synthetic spectra in their pipelines, which help mitigate the scarcity of VMP dwarfs in the training data.
IV.3 MDF Slope Tells Early Assembly Tale
We used the GCE model outlined in Sec. III from our first paper to explore the physical conditions behind different slopes of the metal-poor tail of MDF. The GCE models presented here are pre-selected in Paper I based on the following criteria: 1) [Fe/H] should reach $1.3\pm 0.05$ dex after one Gyr; 2) [Mg/Fe] should reach $0.26\pm 0.05$ dex after one Gyr; 3) [Mg/Fe] increases above 0.3 dex with a minimum increase of 0.05 dex while [Fe/H] surpasses -0.9 dex in the next Gyr.
We assigned each GCE model to one of three slope categories, $k=0.6$, $1.0$, or $1.4$, by comparing its generated MDF to three reference MDFs representing these slope classes. For each model, we first computed a normalized metallicity histogram over the common $\mathrm{[Fe/H]}$ bins, and then evaluated a chi-square-like mismatch statistic with respect to each reference MDF,
$\chi^{2}=\sum_{i}\frac{(M_{i}-R_{i})^{2}}{R_{i}+10^{-10}},$ (4)
where $M_{i}$ is the model MDF in bin $i$ and $R_{i}$ is the corresponding reference MDF. Each model was assigned to the category that gave the smallest mismatch. We retained only models with a minimum mismatch below a threshold of $\chi^{2}<1.5$; models with larger deviations were discarded. This procedure therefore grouped models according to which reference MDF shape they most closely resembled, rather than by fitting each model slope independently.